a total of 34 feet of 1x4 lumber is used around the perimeter of the gate. if the diagonal brace is 13 feet long, then what are the length and width of the gate.
Answer by ewatrrr(24785)

 
Hi,
.
L + w = 17ft L+=+17-w
(17-w)^2 + w^2 = 13^2 Pythagorean Theorem
2w^2 - 34w +120 = 0
w^2 - 17w + 60 = 0
(w-12)(w-5) = 0
Dimensions: 12' by 5'